On July 1st geret pond was 25ft deep since that date,the water level had dropped 2/3ft per week, for what dates will the water level not exceed 21ft
Answer by jorel1380(3719)

2/3x=25-21=4
2x=12
x=6
The water level will not exceed 21 ft. after 6 weeks..